# PV-S3: Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

### v1

@@ 4.1 Quantitative Evaluation
Regarding the performance on UCF-EL dataset with the semi-supervised learning scenario with 20% labeled data, we record average values for metrics as IoU of 67%, precision of 88.90%, recall of 72.50%, and F1-score of 78.17% whereas with 20% labeled data in the supervised learning yielded mean IoU of 8.76%, precision of 26.27%, recall of 25.79% and F1-score of 14.73%. Based on the results we can observe that semi-supervised learning with PV-S3 has yielded significant improvement in results as compared to fully supervised learning with same amount of labeled dataset. The scores with the baseline model of supervised learning with 20% labeled data and semi-supervised learning with 20% labeled data has been summarized in Table  1. To further evaluate the robustness of PV-S3, we perform three experimental runs with different subsets of labeled samples which are randomly selected from the training dataset. The mean and standard deviation of performance metrics for each defect class are shown in Table  2. We observe that the proposed SSL approach significantly outperform the supervised baseline in all metrics consistently, demonstrating its robustness.
@@
Crack 25.29 89.00 26.66 89.75 83.13 98.00 40.37 93.69
Contact 6.89 49.00 10.49 79.50 16.73 57.00 12.89 66.39
Interconnect 0.85 40.00 2.58 88.81 1.26 43.00 1.69 57.94
Corrosion 2.03 90.00 65.37 97.54 2.05 92.00 3.98 94.68
Average 8.76 67.00 26.27 88.90 25.79 72.50 14.73 78.17
@@
Crack $26.20\pm 0.66$ $89.05\pm 0.38$ $26.66\pm 0.63$ $90.43\pm 0.78$ $83.50\pm 1.01$ $97.97\pm 0.24$ $41.41\pm 0.73$ $94.05\pm 0.33$
Contact $6.95\pm 0.06$ $49.95\pm 1.02$ $11.06\pm 1.55$ $77.16\pm 1.17$ $16.92\pm 0.78$ $58.52\pm 1.71$ $12.67\pm 0.89$ $66.95\pm 0.92$
Interconnect $0.7\pm 0.26$ $41.05\pm 0.85$ $2.13\pm 0.78$ $92.07\pm 2.90$ $1.03\pm 0.38$ $42.86\pm 0.81$ $1.39\pm 0.51$ $58.46\pm 0.55$
Corrosion $2.48\pm 0.61$ $86.09\pm 7.37$ $62.40\pm 4.86$ $92.24\pm 6.99$ $2.02\pm 0.06$ $92.62\pm 2.02$ $3.92\pm 0.13$ $92.33\pm 4.36$
Average $9.06\pm 0.09$ $66.53\pm 1.35$ $26.53\pm 1.43$ $88.17\pm 1.94$ $25.86\pm 0.05$ $72.99\pm 0.35$ $14.85\pm 0.05$ $78.00\pm 0.78$
@@
With 100% labeled data using fully supervised learning, the Deeplabv3+ model achieves mean metrics of IoU: 70.84%, precision: 86.29%, recall: 80.12%, and F1-score: 82.03% (Table  3). Comparing these results with our 20% labeled semi-supervised learning approach, PV-S3 yields comparable results to fully supervised learning, with even higher precision. Leveraging 100% labeled data in the semi-supervised framework with PV-S3 significantly improves performance, yielding mean IoU: 74.75%, precision: 91.30%, recall: 79.96%, and F1-score: 84.24%. This improvement underscores the benefits of a larger dataset in better understanding and segmenting defects in PV module EL images.
@@
Crack 88.38 91.78 90.94 93.96 96.91 97.54 93.83 95.72
Contact 52.92 58.29 80.75 78.70 60.56 69.20 69.21 73.65
Interconnect 59.70 53.89 90.70 94.93 63.59 55.49 74.76 70.04
Corrosion 82.39 95.26 82.78 97.62 99.43 97.61 90.34 97.57
Average 70.84 74.75 86.29 91.30 80.12 79.96 82.03 84.24
@@
Deeplabv3+ (Ours SL) $100$ $63.66$ $75.33$ $79.66$ $76.66$
PV-S3 (Ours-SSL (20%)) $20$ $67.00$ $88.90$ $72.50$ $78.17$
PV-S3 (Ours-SSL (100%)) $100$ $74.75$ $91.30$ $79.96$ $84.24$
@@
Comparisons with existing Fully Supervised learning on UCF-EL dataset, performed by [20], across key metrics are summarized in Table  4. We observe improved performance with merely 20% labeled samples and significant enhancement in IoU, Precision, and F1-Score using all labeled samples. Our approach, utilizing just 20% of labeled samples, yields a notable absolute enhancement, including a 9.7% increase in IoU, a 29.9% rise in Precision, a 12.75% boost in Recall, and a 20.42% improvement in F1-Score compared to the previous state-of-the-art supervised method, reducing annotation expenses by 80%.
@@
CSB-Dataset-S1 Crack $50.00$ $58.00$ $57.00$ $64.00$ $80.00$ $86.00$ $66.00$ $73.00$
Gridline $74.00$ $83.00$ $82.00$ $91.00$ $88.00$ $91.00$ $85.00$ $91.00$
Inactive $50.00$ $50.00$ $68.00$ $71.00$ $66.00$ $62.00$ $67.00$ $66.00$
Average $58.00$ $63.67$ $69.00$ $75.33$ $78.00$ $79.67$ $72.67$ $76.67$
CSB-Dataset-S2 Crack $61.00$ $56.00$ $78.00$ $71.00$ $74.00$ $72.00$ $76.00$ $71.00$
Gridline $78.00$ $84.00$ $87.00$ $92.00$ $88.00$ $91.00$ $87.00$ $91.00$
Inactive $62.00$ $60.00$ $70.00$ $73.00$ $86.00$ $77.00$ $77.00$ $75.00$
Average $67.00$ $66.67$ $78.33$ $78.67$ $82.67$ $80.00$ $80.00$ $79.00$
CSB-Dataset-S3 Crack $55.00$ $60.00$ $84.00$ $73.00$ $62.00$ $76.00$ $71.00$ $75.00$
Gridline $77.00$ $85.00$ $88.00$ $90.00$ $86.00$ $93.00$ $87.00$ $92.00$
Inactive $62.00$ $60.00$ $65.00$ $87.00$ $86.00$ $90.00$ $75.00$ $83.00$
Average $64.67$ $68.33$ $79.00$ $83.33$ $78.00$ $86.33$ $77.67$ $83.33$
@@
On the Combined Solar Benchmark Dataset (CSB-Dataset) we evaluate PV-S3 on all three subsets with three major defect classes Crack, Gridline and Inactive. The results for four key metrics, IoU, Precision, Recall and F1-Score, are summarized in Table  5. We observe competitive performance with PV-S3 using only 20% labeled samples across all metrics when compared with fully supervised method using 100% labeled samples. We also compare the performance of PV-S3 with existing methods on these datasets. This comparison is shown in Table  6. The performance of PV-S3 is compared with various baseline models which include PSPNet, UNet-12 and UNet-25 using both 20% (Our-SSL) and 100% (Our-FSL) labels. We observe that conventional semantic segmentation models like PSPNet and UNet variants which are UNet-12 and UNet-25 performed poorly in terms of the Intersection of Union as compared to Deeplab variants. In terms of recall score, UNet-25 has yielded the highest recall score of 60.66% while PSPNet yielded the lowest score of 43.0%. We observe that PV-S3 outperforms other baseline models in terms of all metric scores. Using merely 20% of labeled images for training we observed an increase of 30% in IoU and a 3% increase in recall with a reduction in 80% annotation cost compared to the state-of-the-art model on CSB-Dataset, further using our model with 100% labeled images an increase of 35.67% in IoU and 4.67% increase in recall was observed with state-of-the-art model on CSB-Dataset.
@@ 4.2 Discussion and Analysis
We further study some key aspects such as effectiveness of Semi Cross-Entropy (SemiCE) loss in addressing class imbalance, impact of variations in the amount of labeled data on our model’s performance, and the confidence level of our model in segmenting defects.
@@
A significant challenge in our framework was detecting interconnect defects, which, despite their small image count proportion, constitute a notable pixel distribution percentage, emphasizing the need for precise detection. We experimented with adjusting the Semi Cross-Entropy (SemiCE) thresholds to improve the detection of interconnect defects. The adaptation of Semi Cross-Entropy (SemiCE) positive and negative thresholds is a key aspect of our method. The positive threshold, when set lower, includes predictions with lesser confidence, potentially identifying more defects but risking false positives. A higher positive threshold, on the other hand, demands greater confidence for predictions to be considered positive, thus enhancing accuracy. The negative threshold, conversely, focuses on confidently negative predictions. Setting an appropriate negative threshold ensures that only highly confident negative predictions are considered, avoiding undue focus on uncertain negatives and maintaining the robustness of the defect detection process.
@@
IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1
0.0, 0.0 89 93 95 94 51 71 65 68 50 65 71 78 92 94 98 96
0.2, 0.4 89 92 96 94 53 78 63 70 64 83 74 78 81 88 91 89
0.3, 0.2 89 91 98 94 52 77 61 68 44 79 50 61 67 96 69 80
0.4, 0.2 91 93 97 95 52 73 64 68 45 81 48 62 76 98 77 86
0.5, 0.5 91 93 98 95 53 75 64 69 46 97 46 63 85 97 87 92
0.6, 0.0 89 90 98 94 49 79 57 66 40 89 43 58 90 97 92 95
@@
In our experiments, we varied both the SemiCE positive and negative thresholds to observe their impact on Interconnect defect class performance. Performance with different thresholds is shown in Table 7, with original results in Table  1 using initial thresholds of 0.6 (positive) and 0 (negative), chosen for their balanced nature. Through experimentation, we aimed to identify threshold values optimizing the metrics while maintaining model consistency, shedding light on the intricate relationship between prediction confidence and model performance. The best threshold values achieved were 0.2 (positive) and 0.6 (negative), yielding improved performance for the Interconnect defect class with an IoU of 64%, precision of 83%, recall of 74%, and F1-Score of 78%.
@@
We evaluated PV-S3’s performance using different proportions of labeled data (5% to 20%) and compared it with fully supervised learning using 100% labeled data. Results in Figure  8 indicate that semi-supervised learning with 5% labeled data initially underperforms compared to fully supervised learning. However, performance notably improves at 10% labeled data and continues to increase gradually up to 20%, approaching scores achieved with 100% labeled data for IoU, Recall, and F1 score. Notably, precision scores for semi-supervised learning even exceed those of fully supervised learning.
@@
Figure 10: Comparison of confusion matrix for supervised and semi-supervised learning: The image on the left shows the confusion matrix for supervised learning with 20% labeled images and the right image shows semi-supervised learning with 20% labeled images. The x-axis represents the true class labels and y axis represents the predicted class labels. The values are shown in the log scale for better visualization.

### v3

@@ 4.1 Quantitative Evaluation
Regarding the performance on UCF-EL dataset with the semi-supervised learning scenario with 20% labeled data, we record average values for metrics as IoU of 67%, precision of 72.50%, recall of 88.90%, and F1-score of 78.17% whereas with 20% labeled data in the supervised learning yielded mean IoU of 8.76%, precision of 25.79%, recall of 26.27% and F1-score of 14.73%. Based on the results we can observe that semi-supervised learning with PV-S3 has yielded significant improvement in results as compared to fully supervised learning with same amount of labeled dataset. The scores with the baseline model of supervised learning with 20% labeled data and semi-supervised learning with 20% labeled data has been summarized in Table  1. To further evaluate the robustness of PV-S3, we perform three experimental runs with different subsets of labeled samples which are randomly selected from the training dataset. The mean and standard deviation of performance metrics for each defect class are shown in Table  2. We observe that the proposed SSL approach significantly outperform the supervised baseline in all metrics consistently, demonstrating its robustness.
@@
Crack 25.29 89.00 83.13 98.00 26.66 89.75 40.37 93.69
Contact 6.89 49.00 16.73 57.00 10.49 79.50 12.89 66.39
Interconnect 0.85 40.00 1.26 43.00 2.58 88.81 1.69 57.94
Corrosion 2.03 90.00 2.05 92.00 65.37 97.54 3.98 94.68
Average 8.76 67.00 25.79 72.50 26.27 88.90 14.73 78.17
@@
Crack $26.20\pm 0.66$ $89.05\pm 0.38$ $83.50\pm 1.01$ $97.97\pm 0.24$ $26.66\pm 0.64$ $90.43\pm 0.78$ $41.41\pm 0.73$ $94.05\pm 0.33$
Contact $6.95\pm 0.06$ $49.95\pm 1.02$ $16.92\pm 0.78$ $58.52\pm 1.71$ $11.06\pm 1.55$ $77.16\pm 1.17$ $12.67\pm 0.89$ $66.95\pm 0.92$
Interconnect $0.7\pm 0.26$ $41.05\pm 0.85$ $1.03\pm 0.38$ $42.86\pm 0.81$ $2.13\pm 0.78$ $92.07\pm 2.90$ $1.39\pm 0.51$ $58.46\pm 0.55$
Corrosion $2.48\pm 0.61$ $86.09\pm 7.37$ $2.02\pm 0.06$ $92.62\pm 2.02$ $62.40\pm 4.86$ $92.24\pm 6.99$ $3.92\pm 0.13$ $92.33\pm 4.36$
Average $9.06\pm 0.09$ $66.53\pm 1.35$ $25.86\pm 0.05$ $72.99\pm 0.35$ $26.53\pm 1.43$ $88.17\pm 1.94$ $14.85\pm 0.05$ $78.00\pm 0.78$
@@
With 100% labeled data using fully supervised learning, the Deeplabv3+ model achieves mean metrics of IoU: 70.84%, precision: 80.12%, recall: 86.29%, and F1-score: 82.03% (Table  3). Comparing these results with our 20% labeled semi-supervised learning approach, PV-S3 yields comparable results to fully supervised learning, with even higher precision. Leveraging 100% labeled data in the semi-supervised framework with PV-S3 significantly improves performance, yielding mean IoU: 74.75%, precision: 79.96%, recall: 91.30%, and F1-score: 84.24%. This improvement underscores the benefits of a larger dataset in better understanding and segmenting defects in PV module EL images.
@@
Crack 88.38 91.78 96.91 97.54 90.94 93.96 93.83 95.72
Contact 52.92 58.29 60.56 69.20 80.75 78.70 69.21 73.65
Interconnect 59.70 53.89 63.59 55.49 90.70 94.93 74.76 70.04
Corrosion 82.39 95.26 99.43 97.61 82.78 97.62 90.34 97.57
Average 70.84 74.75 80.12 79.96 86.29 91.30 82.03 84.24
@@
Deeplabv3+ (Ours SL) $100$ $70.84$ 80.12 $86.29$ $82.03$
PV-S3 (Ours-SSL (20%)) $20$ $67.00$ $72.50$ $88.90$ $78.17$
PV-S3 (Ours-SSL (100%)) $100$ 74.75 $79.96$ 91.30 84.24
@@
Comparisons with existing Fully Supervised learning on UCF-EL dataset, performed by [17], across key metrics are summarized in Table  4. We observe improved performance with merely 20% labeled samples and significant enhancement in IoU, Precision, and F1-Score using all labeled samples. Our approach, utilizing just 20% of labeled samples, yields a notable absolute enhancement, including a 9.7% increase in IoU, a 13.5% rise in Precision, a 29.15% boost in Recall, and a 20.42% improvement in F1-Score compared to the previous state-of-the-art supervised method, reducing annotation expenses by 80%.
@@
CSB-Dataset-S1 Crack $50.00$ $58.00$ $80.00$ $86.00$ $57.00$ $64.00$ $66.00$ $73.00$
Gridline $74.00$ $83.00$ $88.00$ $91.00$ $82.00$ $91.00$ $85.00$ $91.00$
Inactive $50.00$ $50.00$ $66.00$ $62.00$ $68.00$ $71.00$ $67.00$ $66.00$
Average $58.00$ $63.67$ $78.00$ $79.67$ $69.00$ $75.33$ $72.67$ $76.67$
CSB-Dataset-S2 Crack $61.00$ $56.00$ $74.00$ $72.00$ $78.00$ $71.00$ $76.00$ $71.00$
Gridline $78.00$ $84.00$ $88.00$ $91.00$ $87.00$ $92.00$ $87.00$ $91.00$
Inactive $62.00$ $60.00$ $86.00$ $77.00$ $70.00$ $73.00$ $77.00$ $75.00$
Average $67.00$ $66.67$ $82.67$ $80.00$ $78.33$ $78.67$ $80.00$ $79.00$
CSB-Dataset-S3 Crack $55.00$ $60.00$ $62.00$ $76.00$ $84.00$ $73.00$ $71.00$ $75.00$
Gridline $77.00$ $85.00$ $86.00$ $93.00$ $88.00$ $90.00$ $87.00$ $92.00$
Inactive $62.00$ $60.00$ $86.00$ $90.00$ $65.00$ $87.00$ $75.00$ $83.00$
Average $64.67$ $68.33$ $78.00$ $86.33$ $79.00$ $83.33$ $77.67$ $83.33$
@@
On the Combined Solar Benchmark Dataset (CSB-Dataset) we evaluate PV-S3 on all three subsets with three major defect classes Crack, Gridline and Inactive. The results for four key metrics, IoU, Precision, Recall and F1-Score, are summarized in Table  5. We observe competitive performance with PV-S3 using only 20% labeled samples across all metrics when compared with fully supervised method using 100% labeled samples. We also compare the performance of PV-S3 with existing methods on these datasets. This comparison is shown in Table  6. The performance of PV-S3 is compared with various baseline models which include PSPNet, UNet-12 and UNet-25 using both 20% (Our-SSL) and 100% (Our-FSL) labels. We observe that conventional semantic segmentation models like PSPNet and UNet variants which are UNet-12 and UNet-25 performed poorly in terms of the Intersection of Union as compared to Deeplab variants. In terms of recall score, UNet-25 has yielded the highest recall score of 60.66% while PSPNet yielded the lowest score of 43.0%. We observe that PV-S3 outperforms other baseline models in terms of all metric scores. Using merely 20% of labeled images for training we observed an increase of 30% in IoU with a reduction in 80% annotation cost compared to the state-of-the-art model on CSB-Dataset, further using our model with 100% labeled images an increase of 35.67% in IoU and 0.33% increase in recall was observed with state-of-the-art model on CSB-Dataset.
@@ 4.2 Discussion and Analysis
We further study some key aspects such as ablation study of loss function, effectiveness of Semi Cross-Entropy (SemiCE) loss in addressing class imbalance, impact of variations in the amount of labeled data on our model’s performance, and the confidence level of our model in segmenting defects.
Ablation Study for Loss Function
We conducted an ablation study on the UCF-EL dataset to compare the performance of the proposed Semi Cross-Entropy (SemiCE) loss with the standard Mean Squared Error (MSE) loss, as shown in Table 7. While the overall results with MSE loss are comparable to those obtained using the SemiCE loss, we observed significant differences in the detection of certain defect classes, especially those with class imbalances like Interconnect and Contact defects. The MSE loss treats all defect classes uniformly during loss calculation, which can be suboptimal for underrepresented classes. In contrast, the SemiCE loss incorporates positive and negative thresholds, allowing the model to optimize predictions for each defect class differently. This approach enhances learning by giving appropriate emphasis to classes with fewer samples. Our results indicate that using MSE loss led to lower performance on Interconnect and Contact defects due to its inability to effectively handle class imbalance. However, when using the SemiCE-S2 variant of our loss function, we observed improved performance on these defect classes. This demonstrates the effectiveness of the SemiCE loss in addressing class imbalance and enhancing overall detection accuracy.
IoU Precision Recall F1-Score
Defect Class MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-12 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2
Crack 91.00 89.00 89.00 97.35 98.00 96.00 91.27 89.75 92.00 94.21 93.69 94.00
Contact 51.00 49.00 53.00 59.51 57.00 63.00 77.55 79.50 78.00 68.23 66.39 70.00
Interconnect 45.00 40.00 64.00 49.56 43.00 74.00 82.71 88.81 83.00 63.54 57.94 78.00
Corrosion 86.00 90.00 81.00 96.95 92.00 91.00 87.96 94.68 88.00 92.23 97.57 89.00
Average 67.75 67.00 71.75 75.84 72.50 81.00 84.87 88.90 85.25 79.65 78.17 82.75
Table 7: Ablation study of SemiCE loss function with MSE loss function for PV-S3 model: The table shows ablation study conducted for SemiCE loss function for different threshold with Mean Squared Error (MSE) loss function. Two setting representing different positive and negative threshold is taken, SemiCE-S1 represents the first setting with positive threshold of 0.6 and negative threshold of 0 and SemiCE-S2 represents the second setting with positive threshold of 0.2 and negative threshold of 0.4.
@@
Detecting interconnect defects presents a significant challenge due to class imbalance: these defects appear in a small proportion of images but constitute a substantial percentage of the pixel distribution. To effectively address this, we employ the Semi Cross-Entropy (SemiCE) loss function, which is specifically designed to handle class imbalances by adjusting the weighting of positive and negative samples during training. The SemiCE loss function incorporates adaptive thresholds for positive and negative predictions. By setting a lower positive threshold, we include more predictions with lower confidence as positive samples. This approach increases the recall of defect detection by identifying more potential defects but may introduce additional false positives. Conversely, a higher positive threshold requires greater confidence for a prediction to be considered positive, enhancing precision but potentially missing some defects. The negative threshold focuses on confidently negative predictions. Setting an appropriate negative threshold ensures that only predictions with high confidence of being negative contribute to the loss. This prevents the model from being influenced by uncertain negative samples, which could degrade performance. By carefully tuning these positive and negative thresholds, the SemiCE loss function balances the influence of classes during training. This adjustment allows the model to handle class imbalance effectively, improving its ability to detect interconnect defects with higher accuracy and robustness.
In our experiments, we varied both the SemiCE positive and negative thresholds to observe their impact on Interconnect defect class performance. Performance with different thresholds is shown in Table 8, with original results in Table  1 using initial thresholds of 0.6 (positive) and 0 (negative), chosen for their balanced nature. Through experimentation, we aimed to identify threshold values optimizing the metrics while maintaining model consistency, shedding light on the intricate relationship between prediction confidence and model performance. The best threshold values achieved were 0.2 (positive) and 0.4 (negative), yielding improved performance for the Interconnect defect class with an IoU of 64%, precision of 74%, recall of 83%, and F1-Score of 78%.
@@
Pos, Neg IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1
0.0, 0.0 89 95 93 94 51 65 71 68 50 71 65 78 92 98 94 96
0.2, 0.4 89 96 92 94 53 63 78 70 64 74 83 78 81 91 88 89
0.3, 0.2 89 98 91 94 52 61 77 68 44 50 79 61 67 69 96 80
0.4, 0.2 91 97 93 95 52 64 73 68 45 48 81 62 76 77 98 86
0.5, 0.5 91 98 93 95 53 64 75 69 46 46 97 63 85 87 97 92
0.6, 0.0 89 98 90 94 49 57 79 66 40 43 89 58 90 92 97 95
@@
We evaluated PV-S3’s performance using different proportions of labeled data (5% to 20%) and compared it with fully supervised learning using 100% labeled data. Results in Figure  8 indicate that semi-supervised learning with 5% labeled data initially underperforms compared to fully supervised learning. However, performance notably improves at 10% labeled data and continues to increase gradually up to 20%, approaching scores achieved with 100% labeled data for IoU, Recall, and F1 score. Notably, recall scores for semi-supervised learning even exceed those of fully supervised learning. Further we observe that the recall score for the semi-supervised model trained with 20% labeled data and 80% unlabeled data exceeds that of the model trained with 100% labeled data. Usually, semi-supervised learning methods which utilized unlabelled samples outperform fully supervised methods [48, 39, 49], This counterintuitive result can be attributed to the benefits of semi-supervised learning in leveraging unlabeled data to improve model generalization. The inclusion of unlabeled data acts as a regularizer, preventing the model from overfitting to the labeled data, which might contain noise or biases. In contrast, training on 100% labeled data without the additional unlabeled data may lead to overfitting, especially if the labeled data is limited in diversity.By incorporating a large amount of unlabeled data, the semi-supervised model learns more robust feature representations, capturing the underlying structure of the data more effectively. The semi-supervised approach allows the model to generalize better to unseen data, resulting in higher recall by correctly identifying more true positive instances of defects.
@@
Figure 10: Comparison of confusion matrix for supervised and semi-supervised learning: The image on the left shows the confusion matrix for supervised learning with 20% labeled images and the right image shows semi-supervised learning with 20% labeled images. The x-axis represents the predicted class labels and y axis represents the true class labels. The values are shown in the log scale for better visualization.
@@
Practical Implications and Real-World Applications
The proposed semi-supervised learning method using a mean teacher network demonstrates significant potential for real-world applications in the field of photovoltaics. By effectively detecting defects in photovoltaic (PV) cells using only 20% of labeled data, our approach addresses a critical challenge in the industry: the scarcity and cost of obtaining fully annotated datasets. This reduction in the requirement for labeled data not only lowers the barrier for implementing advanced defect detection systems but also accelerates the deployment of quality control processes in PV manufacturing. The ability to accurately identify defects with limited labeled data can lead to more efficient production lines, early detection of faults, and overall improvement in the reliability and performance of PV systems. Furthermore, the robustness of our method to class imbalances and its adaptability to diverse datasets enhance its applicability across different types of PV cells and manufacturing conditions. By facilitating more accessible and cost-effective defect detection, our research has the potential to significantly impact the quality assurance processes in the photovoltaics industry, contributing to the advancement of renewable energy technologies.